OpenStudy (anonymous):

Which is the maximum value for the vertices for the following: y≤2x−3y≤−13x+4 x≤9 f( x,y )=3x−2y A. -3 B. 3 C. 16 D. 25

OpenStudy (aum):

Can you post a screenshot of the question?

OpenStudy (anonymous):

http://prntscr.com/4h46sb

OpenStudy (aum):

Much better. Plot the following three lines: y = 2x - 3 y = -1/3 * x + 4 x = 9 and find all the intersection points first.

OpenStudy (aum):

If you want to skip graphing, you can find the intersection of the lines: y = 2x - 3 and y = -1/3 * x + 4 y = -1/3 * x + 4 and x = 9 y = 2x - 3 and x = 9 Let me know when you find the three points of intersection.